# Both Ends Count! Just How Good are LLM Agents at Text-to-“Big SQL”?

### v1

@@ 2.1. Limitations of Current Metrics
The major limitation of these metrics is their reliance on all-or-nothing correctness [^2] The corresponding formulas are provided in the Appendix.. In practice, however, a generated query may deviate from the expected result without being entirely invalid, provided that end users may still be able to determine its correctness through simple validation, for example, by detecting an additional column in the projected output. We summarize the possible outcomes of an SQL execution as follows:
@@ 2.2. Proposed Metrics
[…]
@@
where $T_{gold}$ denotes the execution time of the golden query, and $\mathds{1}(V_{i},\hat{V}_{i})$ is an indicator function that tells whether the output of the generated query matches the expected output ( A and B):
(2) $\mathds{1}(V,\hat{V})=\begin{cases}1,&\text{if $\hat{V}$ is contained in expected output $V$}\\ 0,&\text{otherwise}\end{cases},$
@@ 4.2. Big SQL Metrics Zoom In on Performance
In the context of Big Data, text-to-SQL metrics fail to inform model selection. Instead, a Big SQL lens differentiates LLMs more clearly, providing sharper selection criteria. Table 2 shows normalized VES and VES* with respect to the best scoring LLM setup: GPT-4o. As seen in the table, VES* better discriminates accurate models (809.09% dispersion range vs. 54.93% in VES).
VES considers only query execution time and result accuracy (hong2025database-interfaces). However, modern LLMs excel at text-to-SQL: their generated queries often functionally resemble the gold query (even if not identical) yielding similar execution times. At high accuracy levels, this hinders model discrimination, underscoring the role of precision and agent execution time.
@@
A high VES* effectively reflects both accurate and low latency models, ranking GPT-4o highest (Table 2). It also reveals nuances, such as Opus models’ gains from perfect accuracy despite elevated execution times. Normalizing execution times to the fastest model (GPT-4o) shows it leads across all stages, suggesting a clear split between “fast” and “slow” models.
@@
VCES matches the dispersion range of VES* while incorporating cost, as it factors in per-token billing and the expenses of suboptimal query execution. Table 3 presents VCES per LLM and identifies Gemini 3 Flash as the most cost-efficient option due to its low per-token pricing3330.5/3.0/3.0 per input/output token for Gemini 3 Flash (GoogleGeminiPricing2026), vs. 2.5/10.0/10.0 for GPT-4o (OpenAIPricing2026).. Ultimately, VCES complements VES* by enabling the selection of cost-efficient configurations, thereby addressing a significant gap in text-to-SQL metrics that directly ignore cost. We attach CVQ to better support our finding: GPT-4o more than doubles the per-query cost of Gemini 3 Flash due to its lower accuracy, which leads to more failed queries and higher token consumption.
As with execution time results, the leading cost performer in Table 3 dominates across nearly all stages. Mirroring the findings from VES*, there could be a separation of “cheap” and “expensive” models; the former may be particularly effective for stages with less stringent accuracy or latency requirements.
@@ 4.3. The Aftermath of Data Scale
Text-to-SQL metrics do not provide a representative picture of data scale. However, data scale plays a crucial role in text-to-Big SQL, as illustrated in Figure 1. Two key observations emerge. First, as the scale increases, the relative importance of agent performance decreases, while SQL execution becomes more significant. Second, query failures at scale factors (SF) 10 and 1000 should be clearly distinguished; for example, executing an invalid Query 21 from TPC-H on the same cluster is $13.30\times$ more costly at scale factor 1000 than at scale factor 10.
@@
To investigate these concerns, we selected four TPC-H queries that LLMs struggle with (see Figure 2): complex ones (17, 18, 21) with nested subqueries, plus simpler Query 1 on a single table. We used TPC-H business questions as NL inputs and official SQL as ground truth (see Appendix for selection details). Tests ran on Amazon EMR (aws_emr_2026) with r5b.xlarge master/core nodes, 32 core nodes (4 vCPUs each), and 128 GB gp3 EBS volumes.
Figure 2. Average EX across three models for TPC-H queries 1, 17, 18, and 21.
Text-to-SQL metrics data scale, as illustrated in Figure 3. VES (which incorporates SQL execution time) yields constant relationships between models across scale factors. In contrast, SVQ better captures each model’s potential cost loss at varying scales: less accurate models (e.g., Gemini 3 Pro here) pose greater risk at large scale factors, where query errors incur substantially higher costs.
Consequently, failing a query at SF 1000 is far more expensive than at SF 10. Even a modest accuracy gap (such as the 10% difference between Opus 4.5 and GPT-5.2) becomes critically amplified at higher scales.
@@
Our evaluation clarifies the interpretation of Text-to-Big SQL metrics. VES* and VCES provide practical assessments at fixed scale factors, as they normalize generated SQL execution times and costs relative to the ground truth. In contrast, SVQ serves as an essential complement by quantifying the amplified impact of query inaccuracies as data scales.
@@ 6. Conclusion
In this work, we address the surprisingly underexplored domain of “Text-to-Big SQL”: the integration of Text-to-SQL systems within Big Data workflows. We prove the shortcomings of Text-to-SQL benchmarking in this context and we propose tailored Text-to-Big SQL metrics that capture execution efficiency, cost and scaling effects.
We evaluate state-of-the-art production LLMs to demonstrate that our metrics provide an effective framework for assessing Text-to-Big SQL performance and cost. Our contributions pave the way for promising real-world challenges and highlight avenues for integrating our findings into future research endeavors within Text-to-Big SQL.
@@ Acknowledgements.
Appendix A Related Work
@@
Appendix B Text-to-SQL formulas
@@ B.2. Execution Accuracy (EA)
B.3. Valid Efficiency Score
@@
Appendix C BIRD evaluation
C.1. Model Selection
@@ C.2. Detailed Results
Table 4. Per-query BIRD metrics by model (mean $\pm$ std across runs). VES, VES*, VCES and SVQ follow the paper definitions.
@@
Appendix D TPC-H evaluation
D.1. Model and Query Selection
@@
Appendix E Detailed Results
@@
Table 6. TPC-H metrics aggregated by scale factor (mean $\pm$ std across models and queries).
Scale Factor VES VES* VCES SVQ
SF10 0.1701 $\pm$ 0.0689 0.0731 $\pm$ 0.0264 0.7768 $\pm$ 0.2230 0.2679 $\pm$ 0.0992
SF100 0.0894 $\pm$ 0.0368 0.0521 $\pm$ 0.0198 0.3488 $\pm$ 0.0794 0.3931 $\pm$ 0.1484
SF1000 0.0158 $\pm$ 0.0045 0.0139 $\pm$ 0.0041 0.0244 $\pm$ 0.0058 1.7746 $\pm$ 0.6972

### v2

@@ 2.1. Limitations of Current Metrics
The major limitation of these metrics is their reliance on all-or-nothing correctness [^2] The corresponding formulas are provided in the Appendix. And extended version of this work is available on arXiv: https://arxiv.org/abs/2602.21480. In practice, however, a generated query may deviate from the expected result without being entirely invalid, provided that end users may still be able to determine its correctness through simple validation, for example, by detecting an additional column in the projected output. We summarize the possible outcomes of an SQL execution as follows:
@@ 2.2. Proposed Metrics
To encode our notion of query correctness in a measurable metric, we extend the standard text-to-SQL VES metric to account for superfluous columns. Specifically, to quantify the overhead introduced by including irrelevant columns, we compute the column-level precision: $P(S,\hat{S})=\frac{S\cap\hat{S}|}{|\hat{S}|}$, where $S$ is the set of ground-truth columns and $\hat{S}$ is the set of columns in the execution result. This captures the fraction of retrieved columns that are actually relevant, penalizing additional, unnecessary columns without discarding partially correct results ( C).
@@
where $T_{gold}$ denotes the execution time of the golden query, and $\mathds{1}(V_{i},\hat{V}_{i})$ is an indicator function that tells whether the output of the generated query $\hat{V}_{i}$ matches the expected result $V_{i}$ ( A and B):
(2) $\mathds{1}(V,\hat{V})=\begin{cases}1,&\text{if $V$ is contained in expected output $\hat{V}$}\\ 0,&\text{otherwise}\end{cases},$
@@ 4.2. Big SQL Metrics Zoom In on Performance
In the context of Big Data, text-to-SQL metrics fail to inform model selection. Instead, a Big SQL lens differentiates LLMs more clearly, providing sharper selection criteria. Table 2 shows normalized VES and VES* with respect to the best scoring LLM setup: GPT-4o. As shown in the table, the VES* metric better discriminates accurate models (809.09% dispersion range versus 54.93% in VES). The main reason is that VES considers only query execution time and result accuracy (hong2025database-interfaces). However, modern LLMs excel at text-to-SQL: their generated queries often functionally resemble the gold query (even if not identical) yielding similar execution times. At high accuracy levels, this hinders model discrimination. Our proposed metric addresses this limitation by breaking ties, favoring LLMs that combine efficient agentic interaction with minimal projection overhead (i.e., avoiding superfluous columns).
@@
A high VES* effectively reflects both accurate and low latency models, ranking GPT-4o highest (Table 2). It also reveals nuances, such as the Opus models benefiting from perfect accuracy despite incurring higher execution times. When normalizing execution times relative to the fastest model (GPT-4o), GPT-4o consistently “wins” across all stages, indicating a clear separation between“fast” and “slow” model classes. VES is unable to capture this system behavior.
@@
VCES matches the discriminative power of VES* while incorporating cost, as it factors in per-token billing and the expenses of suboptimal query execution. Table 3 presents VCES per LLM and identifies Gemini 3 Flash as the most cost-efficient option due to its low per-token pricing3330.5/3.0/3.0 per input/output token for Gemini 3 Flash (GoogleGeminiPricing2026), vs. 2.5/10.0/10.0 for GPT-4o (OpenAIPricing2026).. Ultimately, VCES complements VES* by enabling the selection of cost-efficient LLM setups, thus addressing a significant gap in text-to-SQL metrics that directly ignore cost. We attach CVQ to better support our finding: GPT-4o more than doubles the per-query cost of Gemini 3 Flash due to its lower accuracy, which leads to more failed queries and higher token consumption.
As with execution time results, the leading cost performer in Table 3 dominates across nearly all stages. Interestingly, the results reveal a clear trade-off between latency-optimal and cost-efficient models: for example, GPT-4o ranks first in execution speed but achieves only half the cost-efficiency of Gemini 3 Flash, which is slightly slower but much cheaper. This distinction suggests that cost-efficient models may be preferable for stages with relaxed latency requirements, while VES*-optimal models are ideal when latency is critical.
@@ 4.3. The Aftermath of Data Scale
[…]
@@
The remaining key question is the extent to which a new metric is needed to capture data scaling effects. To examine this, we chose four TPC-H queries where LLMs deliver poor accuracy (Figure 2): three complex ones (Q17, Q18, Q21) with nested subqueries, plus one simple single-table query (Q1). We used TPC-H business questions as NL inputs and official SQL as ground truth (see Appendix for details). Tests ran on Amazon EMR (aws_emr_2026) with r5b.xlarge master/core nodes, 32 core nodes (4 vCPUs each), and 128 GB gp3 EBS volumes.
Figure 2. Execution accuracy across three models for TPC-H queries 1, 17, 18, and 21. Accuracy is computed as the average value of Equation 2 over 10 zero-shot translations of each corresponding natural language query.
As illustrated in Figure 3. VES (which incorporates SQL execution time) yields constant relationships between models across scale factors. In contrast, CVQ better captures each model’s potential cost loss at varying scales: less accurate models (e.g., Gemini 3 Pro here) pose greater risk at large scale factors, where query errors incur substantially higher costs. Consequently, failing a query at SF 1000 is far more expensive than at SF 10. Even a modest accuracy gap (such as the 10% difference between Opus 4.5 and GPT-5.2) becomes critically amplified at higher scales.
@@
Our evaluation clarifies the interpretation of text-to-Big SQL metrics. VES* and VCES provide practical assessments at fixed data scale factors, as they normalize generated SQL execution time and cost to the ground truth. In contrast, CVQ serves as an essential complement by quantifying the amplified impact of query inaccuracies as data scales.
@@ 6. Conclusion
In this work, we have started addressing the surprisingly underexplored domain of “text-to-Big SQL”: the integration of text-to-SQL systems within Big Data engines. We have shown that standard text-to-SQL benchmarks fall short and, for the first time, introduced text-to-Big SQL measures that reflect the true interaction between agents, LLMs, Big Data systems, and data scale. We have evaluated state-of-the-art production LLMs to demonstrate that our metrics provide an effective assessment framework for text-to-Big SQL. Our work exposes new real-world challenges and guides future research in text-to-Big SQL.
@@ Acknowledgements.
Appendix A Agent Architecture
Figure 5. The architecture of our agent is centered on an executor built with LangGraph. For our evaluation, both the controller and the checker operate on the same LLM. The Spark Session can be connected to either a local deployment or a distributed cluster.
Appendix B Related Work
@@
Appendix C Text-to-SQL formulas
@@ C.2. Execution Accuracy (EA)
C.3. Valid Efficiency Score
@@
Appendix D BIRD evaluation
D.1. Model Selection
@@ D.2. Detailed Results
Table 4. Per-query BIRD metrics by model (mean $\pm$ std across runs). VES, VES*, and VCES follow the paper definitions; SVQ reports the expected cost per valid query.
@@
Appendix E TPC-H evaluation
E.1. Model and Query Selection
@@
Appendix F Detailed Results
@@
Table 6. TPC-H metrics by query, model, and scale factor (mean $\pm$ std across runs).
Query ID Scale Factor Model VES VES* VCES SVQ
1 10 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.5668 $\pm$ 0.0000 0.2133 $\pm$ 0.0064 1.8351 $\pm$ 0.0783 0.1163 $\pm$ 0.0015
GPT-5.2 0.3235 $\pm$ 0.2641 0.1857 $\pm$ 0.1517 3.0724 $\pm$ 2.5113 0.1007 $\pm$ 0.0000
100 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.2654 $\pm$ 0.0000 0.1494 $\pm$ 0.0031 0.8809 $\pm$ 0.0259 0.1697 $\pm$ 0.0015
GPT-5.2 0.1599 $\pm$ 0.1306 0.1170 $\pm$ 0.0955 1.0526 $\pm$ 0.8598 0.1852 $\pm$ 0.0000
1000 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.0484 $\pm$ 0.0000 0.0424 $\pm$ 0.0002 0.0683 $\pm$ 0.0006 0.6199 $\pm$ 0.0015
GPT-5.2 0.0289 $\pm$ 0.0236 0.0271 $\pm$ 0.0221 0.0478 $\pm$ 0.0390 0.9439 $\pm$ 0.0000
17 10 Gemini 3 Pro 0.0519 $\pm$ 0.1556 0.0121 $\pm$ 0.0362 0.1148 $\pm$ 0.3444 0.8777 $\pm$ 0.0000
Opus 4.5 0.1040 $\pm$ 0.2079 0.0284 $\pm$ 0.0570 0.1814 $\pm$ 0.3637 0.6704 $\pm$ 0.0026
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.1025 $\pm$ 0.0004
100 Gemini 3 Pro 0.0332 $\pm$ 0.0997 0.0107 $\pm$ 0.0320 0.0807 $\pm$ 0.2422 1.1488 $\pm$ 0.0000
Opus 4.5 0.0669 $\pm$ 0.1337 0.0247 $\pm$ 0.0494 0.1335 $\pm$ 0.2674 0.8111 $\pm$ 0.0026
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.1357 $\pm$ 0.0004
1000 Gemini 3 Pro 0.0051 $\pm$ 0.0154 0.0039 $\pm$ 0.0116 0.0071 $\pm$ 0.0213 5.2761 $\pm$ 0.0000
Opus 4.5 0.0094 $\pm$ 0.0187 0.0075 $\pm$ 0.0151 0.0113 $\pm$ 0.0225 3.2374 $\pm$ 0.0026
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.7089 $\pm$ 0.0004
18 10 Gemini 3 Pro 0.5006 $\pm$ 0.0000 0.1879 $\pm$ 0.0142 1.8557 $\pm$ 0.1625 0.1014 $\pm$ 0.0019
Opus 4.5 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
GPT-5.2 0.1017 $\pm$ 0.3052 0.0517 $\pm$ 0.1551 1.0504 $\pm$ 3.1511 0.1172 $\pm$ 0.0029
100 Gemini 3 Pro 0.2563 $\pm$ 0.0000 0.1382 $\pm$ 0.0080 0.8541 $\pm$ 0.0561 0.1619 $\pm$ 0.0019
Opus 4.5 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
GPT-5.2 0.0616 $\pm$ 0.1849 0.0389 $\pm$ 0.1166 0.5274 $\pm$ 1.5823 0.1783 $\pm$ 0.0029
1000 Gemini 3 Pro 0.0481 $\pm$ 0.0000 0.0414 $\pm$ 0.0008 0.0592 $\pm$ 0.0012 0.6997 $\pm$ 0.0019
Opus 4.5 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
GPT-5.2 0.0187 $\pm$ 0.0562 0.0159 $\pm$ 0.0477 0.0737 $\pm$ 0.2211 0.5334 $\pm$ 0.0029
21 10 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.3922 $\pm$ 0.1961 0.1984 $\pm$ 0.0992 1.2211 $\pm$ 0.6107 0.2233 $\pm$ 0.0003
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.4713 $\pm$ 0.0000
100 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.2300 $\pm$ 0.1150 0.1462 $\pm$ 0.0731 0.6527 $\pm$ 0.3264 0.3002 $\pm$ 0.0003
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.7582 $\pm$ 0.0000
1000 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.0308 $\pm$ 0.0154 0.0286 $\pm$ 0.0143 0.0241 $\pm$ 0.0121 1.5039 $\pm$ 0.0003
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 4.4840 $\pm$ 0.0000